# Stochastic Optimization with Optimal Importance Sampling

### v2

@@ Abstract
Importance Sampling (IS) is a widely used variance reduction technique for enhancing the efficiency of Monte Carlo methods, particularly in rare-event simulation and related applications. Despite its effectiveness, the performance of IS is highly sensitive to the choice of the proposal distribution and often requires stochastic calibration. While the design and analysis of IS have been extensively studied in estimation settings, applying IS within stochastic optimization introduces a fundamental challenge: the decision variable and the importance sampling distribution are mutually dependent, creating a circular optimization structure. This interdependence complicates both convergence analysis and variance control. We consider convex stochastic optimization problems with linear constraints and propose a single-loop stochastic approximation algorithm, based on a joint variant of Nesterov’s dual averaging, that jointly updates the decision variable and the importance sampling distribution, without time-scale separation or nested optimization. The method is globally convergent and achieves minimal asymptotic variance among stochastic gradient schemes, matching the performance of an oracle sampler adapted to the optimal solution.
@@ 1 Introduction
Importance sampling (IS) is a powerful variance reduction technique that improves the efficiency of Monte Carlo methods. The key idea is to generate input samples from a deliberately distorted distribution, and then reweight the outputs using a likelihood ratio that accounts for the discrepancy between the original and distorted distributions. With a carefully chosen IS distribution, this approach can yield dramatic reductions in variance, particularly in rare-event simulation, often outperforming naive Monte Carlo by several orders of magnitude. For foundational references, see [glynn1989importance, bucklew2004introduction].
@@
Despite its popularity, IS is often described as a “double-edged sword.” Its performance depends critically on the choice of the proposal distribution, which is typically sensitive to the underlying model. While a well-chosen IS distribution can lead to dramatic variance reductions, a poor choice can result in catastrophic performance degradation [owen2000safe, hesterberg1995weighted]. A large body of work has therefore focused on how to design and calibrate effective IS distributions, using either analytical approaches based on large deviations theory [sadowsky1990large, deo2023achieving] or adaptive numerical methods that update the IS parameters based on streaming Monte Carlo samples [egloff2010quantile, bugallo2017adaptive]. The first approach, originating with [siegmund1976importance], leverages the change of measure used to establish a large deviation lower bound [dembo2016lecture] as a principled and often near-optimal importance sampling distribution. However, it typically applies only to analytically tractable models that admit a large deviation principle [sadowsky1990large]. The second, more widely applicable approach is a statistical one, trading analytical precision for flexibility, requiring minimal a priori model knowledge but generally offering weaker theoretical guarantees. In this work, we adopt the latter perspective, focusing on methods that are broadly applicable and emphasize practical versatility.
@@
Our main focus is the use of importance sampling (IS) to solve stochastic optimization problems. Specifically, we consider problems of the form $\min_{\theta\in\Theta}\mathbb{E}_{X\sim\mathbb{P}}\left[F(\theta,X)\right]$, where $\theta$ is a decision variable in a feasible region $\Theta$ and $X$ is a random variable drawn from a distribution $\mathbb{P}$. In settings where the loss function $F(\theta,X)$ involves, for example, indicators of rare events, standard sampling-based optimization algorithms become inefficient due to the scarcity of informative samples. In such cases, IS can significantly improve sample efficiency. However, while there is a substantial literature on IS for estimating probabilities and expectations (see, e.g., [glynn1989importance, owen2000safe, tokdar2010importance] and references therein), the integration of IS into decision-making and stochastic optimization has only recently begun to receive focused attention (see Section 1.2 on related work). Our work contributes to this emerging line of research. In particular, we present what is, to our knowledge, the first framework that bridges state-of-the-art techniques from stochastic first-order optimization and IS parameter calibration.
@@
We now describe our problem setting and its underlying challenges more precisely. We consider a stochastic optimization problem where the objective takes the form $f(\theta):=\mathbb{E}_{X\sim\mathbb{P}}[F(\theta,X)]$, but the expectation is not available in closed form. Instead, we assume access to samples from either the nominal distribution $\mathbb{P}$ or a parametrized class of IS distributions. Our goal is to leverage this IS family to accelerate the optimization process. This setup presents several layers of difficulty. The first challenge stems from the sensitivity of IS performance to the choice of sampling distribution: an effective IS distribution must be tightly matched to the underlying model. In an optimization context, however, the model is inherently uncertain since the optimal decision $\theta^{\star}$ is not known a priori. This creates a fundamental tension: without a good decision, we cannot identify an effective IS distribution; and without an effective IS distribution, sampling is too inefficient to reliably identify a good decision. This circular dependency, termed the “curse of circularity” by parpas2015importance, has recently been revisited by he2023adaptive, who propose selecting the IS distribution via a deterministic mapping from the current model estimate, relying on structural properties of the problem. In general, however, IS calibration, even for a fixed model, requires an auxiliary sampling-based optimization routine such as the cross-entropy method [deboer2005tutorial] or direct variance minimization [owen2000safe]. Constraints on the decision variable introduce a second layer of difficulty. The optimal IS distribution can change discontinuously with the active constraint set, which undermines the stability of IS-based procedures. As a result, one is effectively required to solve two intertwined optimization problems: one for the decision $\theta$, and one for the IS parameters. This interdependence can multiply the computational cost. The third challenge concerns optimality guarantees. Even if the objective is convex in $\theta$, and the IS calibration problem is well-posed, the interaction between the two leads to a lack of joint convexity, making it nontrivial to establish global convergence or optimality for the combined procedure.
@@ 1.1 Contributions
Given the above discussion, our main contribution is a stochastic approximation framework that integrates importance sampling directly into convex stochastic optimization with linear constraints. The core idea is to jointly update the decision variable and the importance sampling distribution using a single-loop stochastic approximation scheme based on a joint variant of Nesterov’s dual averaging (NDA). The resulting method operates without time-scale separation or nested optimization and applies to broad families of importance sampling distributions. Our main technical contributions are summarized below.
@@
    Global convergence without joint convexity. We establish global convergence of the joint decision and sampling iterates generated by the proposed algorithm under convexity of the objective function and standard regularity conditions on the importance sampling family, including log-convex likelihood ratios as in exponential tilting or mean translation. These guarantees hold despite the fact that the combined optimization problem over the decision and sampling parameters is not jointly convex. Moreover, while the transient behavior and convergence rate of the decision iterates may depend on the chosen importance sampling scheme, we show that their limit is invariant. This invariance implies convergence of the importance sampling parameters to their respective optima.
@@
    Asymptotically optimal variance without prior importance sampling knowledge. We show that the proposed joint scheme resolves the circular dependence between decision optimization and importance sampling calibration in the asymptotic regime. In particular, the averaged decision iterates achieve minimal asymptotic variance among stochastic gradient schemes, matching the performance of an oracle that has access to the optimal importance sampling distribution in advance. Although the result is asymptotic in nature, it establishes that the convergence rate of the decision variable cannot be improved in the limit.
@@ 1.2 Related Work
This work lies at the intersection of two large bodies of literature: stochastic optimization and importance sampling. Given the breadth of each field, we focus our discussion on contributions that directly address their intersection.
@@
The use of importance sampling (IS) in stochastic optimization has primarily been studied in the context of specific objectives, such as quantile, value-at-risk (VaR), and conditional value-at-risk (CVaR) estimation. egloff2010quantile propose a stochastic approximation scheme to asymptotically identify a minimum-variance importance sampler for quantile estimation. While they establish almost sure convergence to the desired quantile, they do not analyze the asymptotic variance of the resulting estimator. pan2020adaptive develop an adaptive IS approach for quantile estimation using a two-layer model, where the inner layer employs a heuristic metamodel. Despite this heuristic structure, their method achieves convergence to the quantile and demonstrates variance reduction relative to standard Monte Carlo. Among the works most closely related to ours are he2023adaptive and Bardou2009computing, which we now discuss in more detail.
@@
he2023adaptive consider adaptive IS for stochastic root-finding problems and demonstrate that a stochastic approximation method, coupled with an IS scheme adapted to the current iterate, can achieve the same asymptotic variance as one adapted to the optimal solution. They refer to this property as minimax optimality. While their work directly tackles the circularity issue in IS calibration, it assumes knowledge of the mapping from decisions to optimal IS parameters, which limits applicability to problems with sufficient analytical structure—typically those where large deviations theory can be used to derive this mapping. In contrast, our method performs stochastic calibration of IS parameters without requiring such mappings to be known in advance. Furthermore, our framework provides global convergence guarantees for general convex objectives with linear constraints, whereas he2023adaptive focus on local convergence in unconstrained, mostly univariate settings.
@@
Bardou2009computing study IS for VaR and CVaR estimation, proposing a stochastic approximation scheme that asymptotically identifies a minimum-variance IS distribution. Under certain convexity conditions on the IS class, they establish a central limit theorem with optimal asymptotic variance. Our work extends this framework in several key directions: (i) we consider general multivariate stochastic optimization problems rather than scalar objectives; (ii) our algorithm handles linear constraints via a recent stochastic approximation technique introduced by duchi2021asymptotic; and (iii) we operate under bounded IS parameter spaces, which simplifies the procedure relative to Bardou2009computing, who employ an auxiliary IS step to handle unbounded domains.
@@
Beyond the IS literature, our work is also connected to broader questions in stochastic optimization. The performance of stochastic algorithms, particularly in modern large-scale settings such as deep learning [bengio2012practical], is often highly sensitive to hyperparameter tuning. This positions our approach within the emerging literature on meta-optimization [hospedales2021meta], where higher-level optimization problems govern the tuning of algorithmic components such as learning rates or sampling distributions.
@@
Finally, we emphasize that while our method guarantees asymptotic optimality within the chosen IS class, we do not make claims about how specific IS families accelerate convergence in practice. This is highly problem-dependent and lies beyond the scope of this work. For recent efforts in this direction, we refer to deo2023achieving, deo2024importance, who study exponential tilting and self-structuring transformations to accelerate CVaR optimization.
@@ 2 Asymptotic Optimality in Stochastic Optimization
Stochastic optimization problems of the form (SO) are ubiquitous in a myriad of domains, including machine learning, stochastic control, queuing theory, portfolio selection and risk management; see schneider2007stochastic, shapiro2021lectures. We consider convex stochastic optimization problems which satisfy the following structural assumptions.
@@
In most practical settings, however, neither the objective function $f:\Theta\to\mathbb{R}$ nor its gradient $\nabla f:\Theta\to\mathbb{R}^{s}$ can be efficiently evaluated, as this would require high-dimensional integration. Instead, an optimizer typically only has access to stochastic gradients $G(\theta,X)$, which represent the gradient of our objective function only in expectation. Optimization algorithms which attempt to solve the problem (SO) using only stochastic gradients are generally known under the name stochastic approximation algorithms, and have been studied since the seminal paper by robbins1951stochastic.
@@ 2.1 Unconstrained Stochastic Optimization
To build intuition, we first consider unconstrained stochastic optimization problems, where $\Theta=\mathbb{R}^{s}$. As the objective function $f$ is continuously differentiable and convex, solving the problem (SO) is equivalent to finding the root of its gradient function [bertsekas2009convex, Proposition 5.4.7], i.e., problem (SO) is equivalent to the first-order optimality condition
@@
with arbitrary $\theta_{0}\in\mathbb{R}^{s}$, and with step-sizes $\alpha_{n}$ satisfying the classical convergence conditions $\sum_{i=1}^{\infty}\alpha_{i}=\infty$ and $\sum_{i=1}^{\infty}\alpha_{i}^{2}<\infty$. The standard choice for the step-sizes in RM-SA is $\alpha_{n}=\alpha/n$ for some constant $\alpha$. Despite its simplicity, one drawback of this method is that its optimal version (having the smallest asymptotic variance [polyak1992acceleration], [asmussen2007stochastic, Chapter VIII-3]; see below in equation (4)), under Assumption 2.1(iv), requires forming $K=\nabla^{2}f(\theta^{\star})^{-1}$, which, in turn, requires knowledge of the unknown optimal solution $\theta^{\star}$. Moreover, it has been observed that the RM-SA method is sensitive to the choice of the step-sizes, often resulting in poor practical performance (see the discussion in [spall2005introduction, Section 4.5.3], and the elucidating example in [nemirovski2009robust, Section 2.1]).
An important improvement of the RM-SA method is the Polyak-Ruppert stochastic approximation (PR-SA) iteration, proposed independently by Polyak [polyak1990new, polyak1992acceleration] and Ruppert [ruppert1988efficient], which takes long step-sizes coupled with a subsequent averaging of the iterates,
@@
for some $\theta_{0}\in\mathbb{R}^{s}$. A standard choice for the step-sizes in PR-SA is $\alpha_{n}=\alpha/n^{\gamma}$, for $\gamma\in(1/2,1)$ and some constant $\alpha$ (see [polyak1992acceleration, Assumption 3.4]). Under mild regularity conditions, in particular Assumption 2.1(iv), polyak1992acceleration prove that $\bar{\theta}_{n}\overset{\text{a.s.}}{\to}\theta^{\star}$ and that the following central limit theorem (CLT) holds:
@@
That is, the convergence of $\bar{\theta}_{n}$ to $\theta^{\star}$ takes place asymptotically at order $\mathcal{O}(1/\sqrt{n})$, while the variance $(\nabla^{2}f(\theta^{\star}))^{-1}\text{Var}_{X\sim\mathbb{P}}\left[G(\theta^{\star},X)\right](\nabla^{2}f(\theta^{\star}))^{-1}$ can be interpreted as a characterization of the convergence speed (at this order). Importantly, the asymptotic variance of the PR-SA method is optimal among the stochastic gradient-based methods (see [asmussen2007stochastic, Chapter 8, Section 4]); that is, the PR-SA iterates enjoy minimal (in a semidefinite sense) asymptotic variance among all estimates based on stochastic gradients.
The distance of the iterates $\bar{\theta}_{n}$ to the root $\theta^{\star}$ is one particular sub-optimality measure. Following the optimality condition (1), another natural sub-optimality measure is the size of the gradients $\nabla{f}(\bar{\theta}_{n})$ [nesterov2012make]. In fact, this optimality criterion has seen a surge of recent interest in the optimization community, since it generalizes more naturally to nonconvex problems [foster2019complexity]. Under the same regularity assumptions as in polyak1992acceleration, and since the gradient function is continuous (by Assumption 2.1(ii)), the Mann–Wald theorem ensures that $\nabla{f}(\bar{\theta}_{n})\overset{\text{a.s.}}{\to}0$. Moreover, using the optimality condition (1) and the fact that the gradient function is continuously differentiable in a neighborhood of $\theta^{\star}$ (by Assumption 2.1(iii)), the delta method can be employed to recover the following CLT:

### v3

@@ Abstract
Importance Sampling (IS) is a widely used variance reduction technique for enhancing the efficiency of Monte Carlo methods, particularly in rare-event simulation and related applications. Despite its effectiveness, the performance of IS is highly sensitive to the choice of the proposal distribution and often requires stochastic calibration. While the design and analysis of IS have been extensively studied in estimation settings, applying IS within stochastic optimization introduces a lesser-known fundamental challenge: the decision variable and the importance sampling distribution are mutually dependent, creating a circular optimization structure. This interdependence complicates both convergence analysis and variance control. In this paper, we consider the generic setting of convex stochastic optimization with linear constraints. We propose a single-loop stochastic approximation algorithm, based on a variant of Nesterov’s dual averaging, that jointly updates the decision variable and the importance sampling distribution, notably without time-scale separation or nested optimization. The method is globally convergent and achieves the minimal asymptotic variance among stochastic gradient schemes, which moreover matches the performance of an oracle sampler adapted to the optimal solution and thus effectively resolves the circular optimization challenge.
@@ 1 Introduction
Importance sampling (IS) is a powerful variance reduction technique that improves the efficiency of Monte Carlo methods. The key idea is to generate input samples from a deliberately distorted distribution, and then reweight the outputs using a likelihood ratio that accounts for the discrepancy between the original and distorted distributions. With a carefully chosen IS distribution, this approach can yield dramatic reductions in variance, particularly in rare-event simulation, often outperforming naïve Monte Carlo methods by several orders of magnitude. For foundational references, see [Bucklew and Bucklew, 2004; Juneja and Shahabuddin, 2006; Blanchet and Lam, 2012].
@@
Despite its popularity, IS is often described as a “double-edged sword”. Its performance depends critically on the choice of the proposal distribution, which is typically sensitive to the underlying model. While a well-chosen IS distribution can lead to dramatic variance reductions, a poor choice can result in catastrophic performance degradation Glynn and Iglehart [1989]; Sadowsky [2002]; L’ecuyer et al. [2010]; Owen and Zhou [2000]; Hesterberg [1995]. A large body of work has therefore focused on how to design and calibrate effective IS distributions, using either analytical approaches based on large deviations theory or adaptive search methods that optimize IS parameters over suitable classes of distributions. The first approach, originating with Siegmund [1976], leverages the change of measure used to establish large deviations rates Dembo [2016] as a principled and often near-optimal importance sampling distribution, and has been applied to various systems across problem domains such as queueing Dupuis et al. [2009]; Blanchet and Lam [2014]; Kroese and Nicola [1999]; Ridder [2009], finance and insurance Glasserman et al. [2008]; Glasserman and Li [2005]; Asmussen and others [1985]; Collamore [2002]; Chen et al. [2019], reliability Rubino and Tuffin [2009]; Heidelberger [1995]; Nicola et al. [2002], biology and multiscale processes Sandmann [2009]; Dupuis et al. [2012]; Vanden-Eijnden and Weare [2012]. However, it typically requires analytically tractable models that admit a large deviations principle sufficiently explicit to be converted into efficient algorithms. The second approach is more versatile, in that it trades analytical precision for flexibility and requires less a priori model knowledge, but generally offers weaker theoretical guarantees. A common approach is the so-called cross-entropy method that casts the IS parameter search problem as a Kullback–Leibler divergence minimization against the oracle best IS distribution De Boer et al. [2005a]; Rubinstein and Kroese [2004]. This typically amounts to a sequential search process involving iterative updates of the IS parameters in order to control the underlying estimation variances. Besides this approach, other effective adaptive methods or enhancements to cross-entropy have been proposed Egloff and Leippold [2010]; Bugallo et al. [2017]; Botev et al. [2013]; Chan and Kroese [2012]; Grace et al. [2014]; Parpas et al. [2015]. In this work, we adopt the latter perspective, focusing on methods that are broadly applicable and emphasize practical versatility.
@@
Specifically, we depart from the probability estimation problem that is extensively studied in the rare-event literature described above. Our main focus is the use of importance sampling (IS) for solving stochastic optimization problems that involve rare events. More precisely, we consider problems of the form $\min_{\theta\in\Theta}\mathbb{E}_{X\sim\mathbb{P}}\left[F(\theta,X)\right]$, where $\theta$ is a decision variable in a feasible region $\Theta$ and $X$ is a random variable drawn from a distribution $\mathbb{P}$. In settings where the loss function $F(\theta,X)$ involves, for example, indicators of rare events, standard sampling-based optimization algorithms become inefficient due to the scarcity of informative samples. In such cases, IS can significantly improve sample efficiency. However, while there is a substantial literature on IS for estimating probabilities and expectations, the integration of IS into decision-making and stochastic optimization has only recently begun to receive focused attention (see Section 1.2 on related work). Our work contributes to this emerging line of research. In particular, we present what is, to the best of our knowledge, the first framework that bridges state-of-the-art techniques from stochastic first-order optimization and IS parameter calibration.
@@
We now describe our problem setting and its underlying challenges more precisely. We consider a stochastic optimization problem where the objective takes the form $f(\theta):=\mathbb{E}_{X\sim\mathbb{P}}[F(\theta,X)]$, but the expectation is not available in closed form. Instead, we assume access to samples from the nominal distribution $\mathbb{P}$ as well as from a parametrized class of IS distributions. Our goal is to leverage this IS family to accelerate the optimization process. This setup presents several layers of difficulty. The first challenge stems from the sensitivity of IS performance to the choice of sampling distribution: an effective IS distribution must be tightly matched to the underlying model. In an optimization context, however, the model is inherently uncertain, since the optimal decision $\theta^{\star}$ is not known a priori. This creates a fundamental tension: without a good decision, we cannot identify an effective IS distribution; and without an effective IS distribution, sampling is too inefficient to reliably identify a good decision. This circular dependency has recently been highlighted and studied by He et al. [2023], who propose selecting the IS distribution through an iterative process. He et al. [2023], however, crucially assumes an a priori known scheme that maps model parameters into effective IS parameters. This important presumption implicitly limits applicability to tractable models that are amenable to analysis via large deviations techniques, in which case a computable mapping can be derived to obtain IS parameters from the model specifications. In general, as we have described earlier, we aim to offer a more versatile and widely applicable approach. In this case, IS calibration, even for a fixed model, requires an auxiliary sampling-based optimization routine such as the cross-entropy method De Boer et al. [2005b], direct variance minimization Owen and Zhou [2000], or their variants. As a first goal of this paper, we develop an efficient algorithm that simultaneously handles the optimization of the IS parameters and the underlying decision problem. With this goal in mind, we study our approach under two additional, substantial, generality considerations from an optimization perspective. First, we allow for constraints on the decision variable that introduce another layer of difficulty. The optimal IS distribution can change discontinuously with the active constraint set, which can undermine the stability of IS-based procedures. As a result, one is effectively required to solve two intertwined optimization problems: one for the decision $\theta$, and one for the IS parameters, and this interdependence can multiply the computational cost. Second, we aim to establish global optimality guarantees. Even if the objective is convex in $\theta$, and the IS calibration problem is well-posed, the interaction between the two leads to a lack of joint convexity, making it nontrivial to establish global convergence or optimality for the combined procedure.
@@ 1.1 Contributions
Given the above discussion, our main contribution is a stochastic approximation framework that integrates versatile, search-based importance sampling directly into convex stochastic optimization with linear constraints. Our core idea is to jointly update the decision variable and the importance sampling distribution using a single-loop stochastic approximation scheme based on a joint variant of Nesterov’s dual averaging (NDA). The resulting method operates without time-scale separation or nested optimization and applies to broad families of importance sampling distributions. Taken together, our work represents, to the best of our knowledge, the first attempt to integrate first-order optimization methods with importance sampling, at a level sophisticated enough to allow the attainment of several notable new guarantees. More precisely, our main technical contributions are summarized below.
@@
    Global convergence without joint convexity. We establish global convergence of the joint decision and sampling iterates generated by the proposed algorithm under convexity of the objective function and standard regularity conditions on the importance sampling family, including log-convex likelihood ratios as in exponential tilting or mean translation. These guarantees hold despite the fact that the combined optimization problem over the decision and sampling parameters is not jointly convex.
@@
    Asymptotically optimal variance without prior importance sampling knowledge. We show that the proposed joint scheme resolves the circular dependence between decision optimization and importance sampling calibration in the asymptotic regime. In particular, the averaged decision iterates achieve minimal asymptotic variance among stochastic gradient schemes, matching the performance of an oracle that has access to the optimal importance sampling distribution in advance. Thus, in the asymptotic sense, the performance of our approach cannot be improved.
@@ 1.2 Related Work
This work lies at the intersection of two large bodies of literature: stochastic optimization and importance sampling. We have already described a range of related work in the introduction. Here, we focus our discussion on contributions that directly address the intersection of these two bodies of literature and that we have not covered yet.
@@
The use of importance sampling (IS) in stochastic optimization has primarily been studied in the context of specific objectives, such as quantile, value-at-risk (VaR), and conditional value-at-risk (CVaR) estimation. Egloff and Leippold [2010] propose a stochastic approximation scheme to asymptotically identify a minimum-variance importance sampler for quantile estimation. While they establish almost sure convergence to the desired quantile, they do not analyze the asymptotic variance of the resulting estimator. Pan et al. [2020] develop an adaptive IS approach for quantile estimation using a two-layer model, where the inner layer employs a heuristic metamodel. Under this structure, their method achieves convergence to the quantile and demonstrates variance reduction relative to standard Monte Carlo, but they do not analyze the magnitude of improvement. Among the works most closely related to ours are He et al. [2023] and Bardou et al. [2009], which we now discuss in more detail in terms of both their connections to and distinctions from our work.
@@
He et al. [2023] consider adaptive IS for stochastic root-finding problems, which can arise from solving optimization problems. They raise the circularity challenge we described in the introduction, where, on one hand, good IS parameters rely on accurate model specifications, including the decision, while, on the other hand, a good decision relies on efficient IS that hinges on good parameters. To tackle this challenge, He et al. [2023] propose an adaptive approach to embed IS into sequential algorithms such as stochastic approximation, and show how appropriate configuration can achieve the same asymptotic variance as if using an IS based on a priori knowledge of the optimal solution. They refer to this property as minimax optimality. While their work, like us, directly tackles the circularity issue in IS calibration, it crucially assumes knowledge of a mapping from decisions to optimal IS parameters, which limits applicability to problems with sufficient analytical structure—typically those where large deviations theory can be used to derive this mapping. In contrast, our method performs stochastic calibration of IS parameters without requiring such mappings to be known in advance. Furthermore, our framework provides global convergence guarantees for general convex objectives with linear constraints, whereas He et al. [2023] focus on local convergence in unconstrained, predominantly univariate settings.
@@
Bardou et al. [2009] study IS for VaR and CVaR estimation, proposing a stochastic approximation scheme that asymptotically identifies a minimum-variance IS distribution. Under certain convexity conditions on the IS class, they establish a central limit theorem with optimal asymptotic variance. Our work extends this framework in several key directions: (i) we consider general multivariate stochastic optimization problems rather than scalar objectives; (ii) our algorithm handles linear constraints via a recent stochastic approximation technique introduced by Duchi and Ruan [2021]; and (iii) we operate under bounded IS parameter spaces, which simplifies the procedure relative to that of Bardou et al. [2009], who employ an auxiliary IS step to handle unbounded domains.
@@
Beyond the IS literature, our work is also connected to broader questions in stochastic optimization. The performance of stochastic algorithms, particularly in modern large-scale settings such as deep learning Bengio [2012], is often highly sensitive to hyperparameter tuning. This positions our approach within the emerging literature on meta-optimization Hospedales et al. [2021], where higher-level optimization problems govern the tuning of algorithmic components such as learning rates or sampling distributions.
@@
Finally, we emphasize that, while our method guarantees asymptotic optimality within the chosen IS class, we do not make claims about how specific IS families accelerate convergence in practice. This is highly problem-dependent and lies beyond the scope of this work. For recent efforts in this direction, we refer to Deo and Murthy [2023, 2024], who study exponential tilting and self-structuring transformations to accelerate CVaR optimization.
@@ 2 Asymptotic Optimality in Stochastic Optimization
Stochastic optimization problems of the form (SO) are ubiquitous in a myriad of domains, including machine learning, stochastic control, queuing theory, portfolio selection and risk management; see Schneider and Kirkpatrick [2007]; Shapiro et al. [2021]. We consider convex stochastic optimization problems which satisfy the following structural assumptions.
@@
In most practical settings, however, neither the objective function $f:\Theta\to\mathbb{R}$ nor its gradient $\nabla f:\Theta\to\mathbb{R}^{s}$ can be efficiently evaluated, as this would require high-dimensional integration. Instead, an optimizer typically only has access to stochastic gradients $G(\theta,X)$, which represent the gradient of our objective function only in expectation. Optimization algorithms which attempt to solve the problem (SO) using only stochastic gradients are generally known under the name stochastic approximation algorithms, and have been studied since the seminal paper by Robbins and Monro [1951].
@@ 2.1 Unconstrained Stochastic Optimization
To build intuition, we first consider unconstrained stochastic optimization problems, where $\Theta=\mathbb{R}^{s}$. As the objective function $f$ is continuously differentiable and convex, solving the problem (SO) is equivalent to finding the root of its gradient function [Bertsekas, 2009, Proposition 5.4.7], i.e., problem (SO) is equivalent to the first-order optimality condition
@@
with arbitrary $\theta_{0}\in\mathbb{R}^{s}$, and with step-sizes $\alpha_{n}$ satisfying the classical convergence conditions $\sum_{i=1}^{\infty}\alpha_{i}=\infty$ and $\sum_{i=1}^{\infty}\alpha_{i}^{2}<\infty$. The standard choice for the step-sizes in RM-SA is $\alpha_{n}=\alpha/n$ for some constant $\alpha$. Despite its simplicity, one drawback of this method is that its optimal version (having the smallest asymptotic variance [Polyak and Juditsky, 1992], [Asmussen and Glynn, 2007, Chapter VIII-3]; see below in equation (4)), under Assumption 2.1(iv), requires forming $K=\nabla^{2}f(\theta^{\star})^{-1}$, which, in turn, requires knowledge of the unknown optimal solution $\theta^{\star}$. Moreover, it has been observed that the RM-SA method is sensitive to the choice of the step-sizes, often resulting in poor practical performance (see the discussion in [Spall, 2005, Section 4.5.3], and the elucidating example in [Nemirovski et al., 2009, Section 2.1]).
An important improvement of the RM-SA method is the Polyak-Ruppert stochastic approximation (PR-SA) iteration, proposed independently by Polyak Polyak [1990]; Polyak and Juditsky [1992] and Ruppert Ruppert [1988], which takes long step-sizes coupled with a subsequent averaging of the iterates,
@@
for some $\theta_{0}\in\mathbb{R}^{s}$. A standard choice for the step-sizes in PR-SA is $\alpha_{n}=\alpha/n^{\gamma}$, for $\gamma\in(1/2,1)$ and some constant $\alpha$ (see [Polyak and Juditsky, 1992, Assumption 3.4]). Under mild regularity conditions, in particular Assumption 2.1(iv), Polyak and Juditsky [1992, Theorem 2] prove that $\bar{\theta}_{n}\overset{\text{a.s.}}{\to}\theta^{\star}$ and that the following central limit theorem (CLT) holds:
@@
That is, the convergence of $\bar{\theta}_{n}$ to $\theta^{\star}$ takes place asymptotically at order $\mathcal{O}(1/\sqrt{n})$, while the variance $(\nabla^{2}f(\theta^{\star}))^{-1}\text{Var}_{X\sim\mathbb{P}}\left[G(\theta^{\star},X)\right](\nabla^{2}f(\theta^{\star}))^{-1}$ can be interpreted as a characterization of the convergence speed (at this order). Importantly, the asymptotic variance of the PR-SA method is optimal among the stochastic gradient-based methods (see [Asmussen and Glynn, 2007, Chapter 8, Section 4]); that is, the PR-SA iterates enjoy minimal (in a semidefinite sense) asymptotic variance among all estimates based on stochastic gradients.
The distance of the iterates $\bar{\theta}_{n}$ to the root $\theta^{\star}$ is one particular sub-optimality measure. Following the optimality condition (1), another natural sub-optimality measure is the size of the gradients $\nabla{f}(\bar{\theta}_{n})$ [Nesterov, 2012]. In fact, this optimality criterion has seen a surge of recent interest in the optimization community, since it generalizes more naturally to nonconvex problems Foster et al. [2019]. Under the same regularity assumptions as in Polyak and Juditsky [1992, Theorem 2], and since the gradient function is continuous (by Assumption 2.1(ii)), the Mann–Wald theorem ensures that $\nabla{f}(\bar{\theta}_{n})\overset{\text{a.s.}}{\to}0$. Moreover, using the optimality condition (1) and the fact that the gradient function is continuously differentiable in a neighborhood of $\theta^{\star}$ (by Assumption 2.1(iii)), the delta method can be employed to recover the following CLT: